# DetectRL: Benchmarking LLM-Generated Text Detection in Real-World Scenarios

### v1

@@ 3.2 In-domain Robustness
Differences in statistical patterns of LLMs pose significant challenges to detector performance.
As illustrated in subsection 3.1, our experiments reveal a notable phenomenon: nearly all zero-shot LLM-generated text detectors exhibit a significant decline in performance when processing texts generated by Claude. This suggests that the effectiveness of detectors is influenced by the specific generative model used, and their performance can deteriorate with varying statistical patterns. We hypothesize that these differences arise from variations in data, architecture, and training methods of the models, though verifying this is difficult due to the opaque nature of black-box models. Moreover, supervised detectors are more affected by the type of generative model than by the domain, particularly in models with larger sizes. For example, Rob-Large achieved an AUROC of only 86.72% and an $F_{1}$ Score of 76.17% on texts generated by Llama-2, whereas X-Rob-Large achieved an AUROC of 91.67% and an $F_{1}$ Score of 82.24% on texts generated by Claude.
@@
As shown in subsection 3.1, our findings indicate that the adversarial perturbation attacks drastically reduce the effectiveness of zero-shot detectors, reducing their performance to an average AUROC of 34.32%, which is less than half compared to their performance under paraphrase attacks. Additionally, data mixing presents a new challenging scenario, resulting in performance levels similar to semantic attacks, with detectors achieving an average AUROC of 52.51%. While prompt attacks, such as few-shot prompts, can generate higher-quality text more aligned with human preferences, their impact on zero-shot detectors is minimal. However, refining LLM-generated texts via human-written prompts still challenges detectors (see Appendix F.1), decreasing their effectiveness by an average of 9.15% AUROC. This finding suggests that prompt-based methods remain a viable means of compromising detector performance. In contrast, supervised detectors consistently maintain robust performance across various attack types, demonstrating their potential for practical applications.
@@ 3.3 Generalization of detectors
As shown in Table 5, our experimental results indicate that detectors trained on less formal stylistic domain data, such as creative writing and social media, exhibit stronger generalization. Their comprehensive performance is 10% better than detectors trained on more formal stylistic domain data, such as academic writing and news writing. The variations in statistical patterns of generative models significantly impact the generalization of detectors. Detectors trained on texts generated by models with similar statistical patterns, such as GPT-3.5, PaLM-2, and Llama-2, generally perform well with each other. However, they struggle with texts generated by Claude. As discussed in Section 3.2, data with perturbation attacks poses the greatest challenge for generalization. Taking LRR as an example, the average AUROC for detectors trained on data with direct prompts, prompt attacks, paraphrase attacks, and data mixing, and then tested on data with perturbation attacks, is only 27.00%. This performance is 75.31%, 30.17%, and 27.62% lower than when tested on data with prompt attacks, paraphrase attacks, and data mixing, respectively. However, detectors trained on data with perturbation attacks do not show a significant decline in performance when tested on other types of attacks. This indicates that perturbation attack data may possess inherent complexities that are particularly challenging to detect.
@@ 3.4 Impact of text length
Similarly, we trained a detector using data from the pivotal length interval and assessed its performance on test sets with varying text lengths. The experimental results, shown in Figure 3 (b), reveal that as test text length increased, the performance of the zero-shot detectors improved steadily. This suggests a positive correlation between zero-shot detectors’ performance and test text length. In contrast, supervised methods showed a rapid performance increase up to the pivotal length interval, followed by a slight decline.
@@ 3.5 Impact of real-world human writing scenarios
We explored a critical question in real-world detection: How do human-driven factors impact detector performance? To investigate this, we simulated various modifications to human-written texts. We introduced paraphrase attacks to mimic text revisions and incorporated spelling errors through perturbation attacks. Moreover, we mixed LLM-generated sentences with human-written content to simulate AI-assisted writing scenarios. Experimental results, as shown in subsection 3.5, indicate that attacks on human-written texts yield markedly different outcomes compared to those on LLM-generated texts. Specifically, paraphrasing attacks on human-written texts effectively confused zero-shot detectors, reducing the AUROC by an average of 4.77%. In contrast, data mixing had a minimal impact on zero-shot detectors’ performance, with only a slight decline of 4.48% in AUROC. This contrasts sharply with the significant 20.29% decline in AUROC when human-written texts were mixed with LLM-generated texts. The resilience of human-written texts to such mixing may be attributed to their inherent complexity, making it difficult for zero-shot detectors to identify the inclusion of LLM-generated content. Interestingly, perturbation attacks on human-written texts appeared to enhance the discernment capabilities of zero-shot detectors, resulting in an average increase of 11.05% in AUROC. Similar trends were observed with supervised detectors. This suggests that human-written texts may inherently contain more adversarial features [41], which are utilized by detectors for identification. Such perturbations can further emphasize these distinctions, leading to improved performance.
@@ A.3 Detection benchmark
Previous work has already dedicated significant effort to the construction of benchmarks for LLM-generated text detection, mainly encompassing early deepfake research such as the TweepFake dataset [62] and the GROVER Dataset [63], as well as prior work on detecting LLM-generated texts like the GPT-2 Output Dataset [^5] https://github.com/openai/gpt-2-output-dataset and TuringBench [8]. HC3 [47] is one of the recent impressive datasets, containing ChatGPT-generated text data in both English and Chinese, covering multi-domain and multi-lingual evaluation. Other benchmarks, such as MGTBench [9], ArguGPT [64], and the DeepfakeTextDetect Dataset [65], also consider texts generated by various LLMs. M4 [66] is a comprehensive dataset recently released, covering multi-domain, multi-lingual, and multi-generator evaluation scenarios.
@@ Appendix C Ethics statement
We developed DetectRL by collecting publicly available human-written texts in high-risk and abuse-pron domains, generating similar texts using advanced and popular LLMs, designing and applying various attack methods for data augmentation. The release of DetectRL aims to advance research on detecting LLM-generated texts, enhancing their robustness and applicability in real-world scenarios. However, while promoting this research, we have also considered the potential for misuse. By making our dataset construction framework publicly available, there’s a possibility that our well-designed attack methodologies could be used to develop defenses that might undermine existing detection systems.
@@ E.2 Detectors settings
Settings $\rightarrow$ Few-Shot Prompt SICO Prompt Muti-models LLM-centered
@@ Appendix F Additional experiment results
For paraphrase attacks, the results in subsection E.2 demonstrate that DIPPER Paraphrase is not an effective tool for assessing the performance of detectors in complex black-box scenarios, as detectors nearly maintain the same identification performance as with direct prompts. However, Polishing using LLMs proves relatively effective, reducing the detector’s AUROC and $F_{1}\text{ Score}$ by approximately 10%, indicating that self-editing texts with LLMs can still diminish detection capabilities. Notably, Back-translation, a widely used paraphrasing method in our daily life, exhibits strong attack capability, significantly reducing the detector’s AUROC and $F_{1}\text{ Score}$ to 28.41% and 10.74%, respectively, a decline of over 40%.
Regarding perturbation attacks, the results in Appendix F underscore their more threatening nature compared to paraphrase attacks. On average, Character-Level, Word-Level, and Sentence-Level Perturbations resulted in about a 40% performance decrease for detectors, which is second only to Back-translation among all the attack methods we evaluated. Interestingly, these three types of perturbations not only showed high consistency but also revealed their potential threat in assessing detector stability. Nonetheless, all supervised detectors still performed well on datasets with the same distribution.
Lastly, we assessed detector performance in data mixing scenarios, as shown in subsection E.2. The results indicate that mixing texts generated by different LLMs is ineffective at confusing detectors. However, LLM-Centered Mixing poses a greater threat, particularly when a quarter of the replacement sample sentences are human-written. This leads to a performance decrease of over 10% for the detectors.

### v3

@@ 3.2 In-domain Robustness
Differences in statistical patterns of LLMs pose significant challenges to detectors.
As illustrated in subsection 3.1, our experiments reveal a notable phenomenon: nearly all zero-shot LLM-generated text detectors exhibit a significant decline in performance when processing texts generated by Claude. This suggests that the effectiveness of detectors is influenced by the type of generative model used to generate the text to be detected, and their performance can deteriorate with varying statistical patterns. We hypothesize that these differences arise from variations in data, architecture, and training methods of the models, though verifying this is difficult due to the opaque nature of black-box models. Moreover, supervised detectors are more affected by the type of generative model than by the domain, particularly in models with larger sizes. For example, Rob-Large achieved an AUROC of only 86.72% and an $F_{1}$ Score of only 76.17% on texts generated by Llama-2, while X-Rob-Large achieved an AUROC of only 91.67% and an $F_{1}$ Score of only 82.24% on texts generated by Claude.
@@
As shown in subsection 3.1, our findings indicate that the adversarial perturbation attacks drastically reduce the effectiveness of zero-shot detectors, reducing their performance to an average AUROC of 38.43%, which is less than half compared to their performance under paraphrase attacks. Additionally, data mixing presents a new challenging scenario, resulting in performance levels similar to paraphrase attacks, with detectors achieving an average AUROC of 59.50%. While prompt attacks, such as few-shot prompting, can generate higher-quality text more aligned with human preferences, their impact on zero-shot detectors is minimal. However, enhancing LLM-generated texts through human-written prompts, such as those used for polishing, continues to pose challenges for detectors (see Appendix F.1), decreasing their effectiveness by an average of 8.97% AUROC. This finding suggests that prompt-based methods remain a viable means of compromising detector performance. In contrast, supervised detectors consistently maintain robust performance across various attack types, demonstrating their potential for practical applications.
@@ 3.3 Generalization of detectors
As shown in Table 5, our experimental results indicate that detectors trained on less formal stylistic domain data, such as creative writing and social media, exhibit stronger generalization. Their comprehensive performance is around 10% AUROC better than detectors trained on more formal stylistic domain data, such as academic writing and news writing. The variations in statistical patterns of generative models significantly impact the generalization of detectors. Detectors trained on texts generated by models with similar statistical patterns, such as GPT-3.5, PaLM-2, and Llama-2, generally perform well with each other. However, they struggle with texts generated by Claude. As discussed in Section 3.2, data with perturbation attacks poses the greatest challenge for generalization. Taking LRR as an example, the average AUROC for detectors trained on data with direct prompts, prompt attacks, paraphrase attacks, and data mixing, and then tested on data with perturbation attacks, is only 27.00%. This performance is 45.31%, 30.17%, and 27.62% lower than when tested on data with prompt attacks, paraphrase attacks, and data mixing, respectively. However, detectors trained on data with perturbation attacks do not show a significant decline in performance when tested on other types of attacks. This indicates that perturbation attack data may possess inherent complexities that are particularly challenging to detect.
@@ 3.4 Impact of text length
Similarly, we trained detectors using data from the pivotal length interval and assessed their performance on test sets with varying text lengths. The experimental results, shown in Figure 3 (b), reveal that as test text length increased, the performance of the zero-shot detectors improved steadily. This suggests a positive correlation between zero-shot detectors’ performance and test text length. In contrast, supervised methods showed a rapid performance increase up to the pivotal length interval, followed by a slight decline.
@@ 3.5 Impact of real-world human writing scenarios
We explored a critical question in real-world detection: How do human-driven factors impact detector performance? To investigate this, we simulated various modifications to human-written texts. We introduced paraphrase attacks to mimic text revisions and incorporated spelling errors through perturbation attacks. Moreover, we mixed LLM-generated sentences with human-written content to simulate AI-assisted writing scenarios. Experimental results, as shown in subsection 3.5, indicate that attacks on human-written texts yield markedly different outcomes compared to those on LLM-generated texts. Specifically, paraphrasing attacks on human-written texts effectively confused zero-shot detectors, reducing the AUROC by an average of 7.95%. In contrast, data mixing had a minimal impact on zero-shot detectors’ performance, with only a slight decline of 3.71% in AUROC. This contrasts sharply with the significant 18.17% decline in AUROC when human-written texts were mixed with LLM-generated texts. The resilience of human-written texts to such mixing may be attributed to their inherent complexity, making it difficult for zero-shot detectors to identify the inclusion of LLM-generated content. Interestingly, perturbation attacks on human-written texts appeared to enhance the discernment capabilities of zero-shot detectors, resulting in an average increase of 10.22% in AUROC. Similar trends were observed with supervised detectors. This suggests that human-written texts may inherently contain more adversarial features [41], which are utilized by detectors for identification. Such perturbations can further emphasize these distinctions, leading to improved performance.
@@ A.3 Detection benchmark
Previous work has already dedicated significant effort to the construction of benchmarks for LLM-generated text detection, mainly encompassing early deepfake research such as the TweepFake dataset [62] and the GROVER Dataset [63], as well as prior work on detecting LLM-generated texts like the GPT-2 Output Dataset [^5] https://github.com/openai/gpt-2-output-dataset and TuringBench [8]. HC3 [47] is one of the recent impressive datasets, containing ChatGPT-generated text data in both English and Chinese, covering multi-domain and multi-lingual evaluation. Other benchmarks, such as MGTBench [9], ArguGPT [64], and the MAGE [11], also consider texts generated by various LLMs. M4 [65] is a comprehensive dataset recently released, covering multi-domain, multi-lingual, and multi-generator evaluation scenarios.
@@ Appendix C Ethics statement
We developed DetectRL by collecting publicly available human-written texts in high-risk and abuse-pron domains, generating similar texts using advanced and popular LLMs, designing and applying various attack methods for data augmentation. The release of DetectRL aims to advance research on detecting LLM-generated texts, enhancing their robustness and applicability of detectors in real-world scenarios. However, while promoting this research, we have also considered the potential for misuse. By making our dataset construction framework publicly available, there’s a possibility that our well-designed attack methodologies could be used to develop defenses that might undermine existing detection systems.
@@ E.2 Detectors settings
Settings $\rightarrow$ Few-Shot Prompt ICO Prompt Muti-models LLM-centered
@@ Appendix F Additional experiment results
For paraphrase attacks, the results in subsection E.2 demonstrate that DIPPER Paraphrase is not an effective tool for assessing the performance of detectors in complex black-box scenarios, as detectors nearly maintain the same identification performance as with direct prompts. However, Polishing using LLMs proves relatively effective, reducing the detector’s AUROC and $F_{1}\text{ Score}$ by approximately 10%, indicating that self-editing texts with LLMs can still diminish detection capabilities. Notably, Back-translation, a widely used paraphrasing method in our daily life, exhibits strong attack capability, significantly reducing the detector’s AUROC and $F_{1}\text{ Score}$ to 30.42% and 12.64%, respectively, a decline of over 40%.
Regarding perturbation attacks, the results in Appendix F underscore their more threatening nature compared to paraphrase attacks. On average, Character-Level, Word-Level, and Sentence-Level Perturbations resulted in a 37.75% AUROC performance decrease for detectors. Among all the attack methods we evaluated, Character-Level Perturbation ranks second only to Back-translation. Interestingly, these three types of perturbations not only showed high consistency but also revealed their potential threat in assessing detector stability. Nonetheless, all supervised detectors still performed well on datasets with the same distribution.
Lastly, we assessed detector performance in data mixing scenarios, as shown in subsection E.2. The results indicate that mixing texts generated by different LLMs is ineffective at confusing detectors. However, LLM-Centered Mixing poses a greater threat, particularly when a quarter of the replacement sample sentences are human-written. This leads to a performance decrease of 13.19% AUROC for the detectors.